# Reaching Curzon-Ahlborn limit in linear response and Whitney limit in nonlinear response in edge mode quantum thermoelectrics and refrigeration

### v2

@@ Abstract
Quantum heat engines and quantum refrigerators are proposed in three-terminal quantum Hall (QH) and quantum spin Hall (QSH) setups with a voltage-temperature probe in both the linear and nonlinear transport regimes. In the linear response regime, we find that efficiency at maximum power approaches the Curzon-Ahlborn limit in both QH and QSH setups. Similarly, in nonlinear response, we find that efficiency at maximum power reaches the Whitney bounds. This is for the first time, we see that in the same setup and using quantum point contacts, the thermoelectric efficiency limits in linear and nonlinear response being achieved.
@@
Introduction: The quantum Hall resistance [1] in two-dimensional electron gas in the presence of a normal uniform magnetic field at low temperature arises as a consequence of chiral edge modes [2], which is famously known as quantum Hall effect. These edge modes are robust against disorder and are immune to backscattering [3] being topologically protected. Complimentary to chiral edge modes, helical edge modes are seen in quantum spin Hall experiments with Mercury Telluride/Cadmium Telluride heterostructures [4, 5, 6], which are also topologically protected. Recently in QH [7, 8, 9, 10, 11] and QSH [12, 13] setups, thermoelectric properties have been investigated using Landaeur-Buttiker formalism [14, 15]. Quantum transport can be described in two regimes. First is linear, wherein the applied voltage and temperature biases are quite small, and thus leads to a linear relation between current and the temperature or voltage biases applied [16]. The second is nonlinear, wherein the applied voltage and temperature biases are not small. In the linear response regime, the efficiency at maximum power ($\eta|_{P_{max}}$) and the maximum efficiency ($\eta_{max}$) are closely related to the figure of merit $Z\theta\hskip 2.84544pt\left(=\frac{GS^{2}\theta}{K}\right)$, where $G$ = conductance, $S$ = Seebeck coefficient, $K$ = thermal conductance and $\theta$ is the reference temperature of the setup. In this regime, there exists a bound to $\eta|_{P_{max}}$, which is given by Curzon-Ahlborn (CA) [17] limit with the efficiency ($\eta|_{P_{max}}$)= $\eta_{c}/2$, where $\eta_{c}=1-\theta_{2}/\theta_{1}$ is the Carnot efficiency with $\theta_{1}$ and $\theta_{2}$ being the temperatures of the hot and cold reservoirs respectively [16]. To achieve this, it is necessary to have a large $Z\theta$. This can be accomplished by ensuring a high Seebeck coefficient ($S$) and low thermal conductance ($K$). Similarly, in the non-linear transport regime, there also exists an upper bound to $\eta|_{P_{max}}$, the Whitney limit [18, 19], i.e., efficiency at maximum power output ($\eta|_{P_{max}^{Wh}}$) = $\eta|_{P_{max}^{Wh}}$ = $\eta_{c}/(1+0.936(1+\theta_{2}/\theta_{1}))$, $N$ being the number of edge modes. This is the nonlinear equivalent of the CA efficiency in linear response and has been derived utilizing BoxCar-type transmission. The study of two-terminal setups in nonlinear response was extended to a three-terminal setup with a voltage probe, and it was shown that the Whitney limits are universal [20]. Similarly, in [21], it has been proved that a generic mesoscopic thermoelectric heat engine can be highly efficient only when the transmission is Box-car type. Boxcar-type scattering can be implemented via a chain of quantum dots [18, 19, 20]. However, electron-electron interaction between quantum dots, which can lead to coulomb blockade effects, is neglected in the calculation in Refs. [18, 19, 20]. This assumption is quite extreme, as in real systems, having a chain of quantum dots without the coulomb blockade effect is almost impossible to implement [16]. In contrast, this work considers quantum point contact (QPC) type scattering, which is easier to implement experimentally [22, 23, 24, 25] than the boxcar-type transmission. In this work, we reach the Whitney bound in the nonlinear transport regime using QPC-type scattering. Previously, in Refs. [18, 19, 20] a QH setup was only considered; however, in this letter, we consider both QH and QSH setups, albeit with the easier-to-implement QPC type contacts for both linear as well as non-linear response regimes.
@@
This letter is divided into two parts. In the first part, we investigate the thermoelectric transport of the QH and QSH setups within the linear response regime and show that efficiency at maximum power approaches Curzon-Ahlborn for both QH and QSH setups. In the second part, we focus on non-linear transport, wherein we see that the efficiency at maximum power approaches the Whitney limit for both QH and QSH setups. The technical details of the calculation are presented in the supplementary material (SM).
Thermoelectricity in QH and QSH setups: We start by discussing the general theory for thermoelectric transport in QH as well as QSH setups within the Landaeur-Buttiker scattering theory in both linear and non-linear regimes [16, 14]. The net charge and heat currents in a multiterminal QH sample, see Sec. II of SM, are given as,
@@
where the transmission probability for an electron to scatter from terminal $\beta$ to terminal $\alpha$ is $T_{\alpha\beta}$. Here, $f_{\alpha(\beta)}=\left(1+e^{(E-\mu_{\alpha(\beta)})/k_{B}\theta_{\alpha(\beta)}}\right)^{-1}$, $\mu_{\alpha(\beta)}$ and $\theta_{\alpha(\beta)}$ are the Fermi function, chemical potential and temperature of terminal $\alpha(\beta)$ respectively.
@@
Similarly, for QSH setup, with helical edge modes, the currents carry an extra spin degree of freedom and are written as $I_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{\alpha}^{\sigma}$ and $J_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{\alpha}^{\sigma}$, where [12],
@@
$\begin{split}I_{\alpha}^{\sigma}&=\frac{e}{h}\int_{-\infty}^{\infty}dE\sum_{\beta}T_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\\ J_{\alpha}^{\sigma}&=\frac{1}{h}\int_{-\infty}^{\infty}dE(E-\mu_{\alpha})\sum_{\beta}T_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\end{split}$ (2)
@@
with the transmission probability for an electron to transmit from terminal $\beta$ with spin $\rho$ to terminal $\alpha$ with spin $\sigma$ being $T_{\alpha\beta}^{\sigma}=\sum_{\rho\in\{\uparrow,\downarrow\}}T_{\alpha\beta}^{\sigma\rho}$. In this three-terminal setup (both QH and QSH) as shown in Fig. 1, $\theta_{1}=\theta+\tau_{1},\theta_{2}=\theta+\tau_{2}$ and $\theta_{3}=\theta+\tau_{3}$, where $\theta$ is the reference temperature and $\tau_{1},\tau_{2}$ and $\tau_{3}$ are the temperature biases applied in terminals 1, 2 and 3 respectively. For our setups, we apply temperature bias only at terminal 1, i.e., $\tau_{1}=\tau$ and $\tau_{2}=0$. We study the thermoelectric properties for a QHE, considering terminal 3 to be a VT probe. We also assume both terminals 1 and 2 to be current probes with $V_{1}=-V$ and $V_{2}=0$. Here, we have considered only one type of energy-dependent transmission via constrictions, which is known as the quantum point contact (QPC) type [28], defined via $T_{\text{QPC}}=[1+\text{exp}(-2\pi(E-E_{l})/\hbar\omega]^{-1}$, where $l$ denotes the constriction index, i.e., $l\in\{1,2\}$. Below a certain threshold energy level $E_{l}$, $T_{QPC}$ vanishes, while above $E_{l}$, it is transparent, while $\hbar\omega$ indicates the width of the transmission resonance, see also Sec. I of SM. Following Refs. [29, 30], a capacitive external gate voltage $V_{g}$ is connected to the QPC constrictions which, in turn, controls the threshold energy $E_{l}$ of QPC, through which we can explore both linear and nonlinear transport regimes for our setups in Fig. 1(a) and (b).
@@
Linear transport in QH and QSH setups: In the linear response regime for the QH case, i.e., when $\mu_{\alpha}(=eV_{\alpha})\ll k_{B}\theta$ and ($\theta_{\alpha}-\theta)\ll\theta$, one can see that the current (both charge and heat) varies linearly with the voltage and temperature biases applied across the sample [16]. Since, terminal 3 is a VT probe i.e., $I_{3}=J_{3}=0$, the charge and heat currents in terminal 1 can be written as,
$\begin{pmatrix}I_{1}\\ J_{1}\end{pmatrix}=\begin{pmatrix}L_{eV}&L_{e\theta}\\ L_{hV}&L_{h\theta}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (3)
@@
where the Onsager matrix elements are $L_{eV},L_{e\theta},L_{hV},L_{h\theta}$, see Eq. (31) in SM. These Onsager matrix elements are used to find the transport coefficients such as conductance ($G$), Seebeck coefficient ($S$), Peltier coefficient ($\Pi$), and the thermal conductance ($K$) and are given as,
$G$ $=L_{eV},\quad S=\frac{L_{e\theta}}{L_{eV}},\quad\Pi=\frac{L_{hV}}{L_{eV}},\quad K=L_{h\theta}-\frac{L_{hV}L_{e\theta}}{L_{eV}}.$ (4)
@@
Using a similar procedure, one can also derive the Onsager matrix elements in the QSH case for charge current $I_{\alpha}^{\sigma}$ and heat current $J_{\alpha}^{\sigma}$ by imposing the VT probe condition ($I_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{3}^{\sigma}=0$ and $J_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{3}^{\sigma}=0$) and they can be written in the same form as Eq. (3) (see, Sec. III.1.1 in SM) as
$\begin{pmatrix}I_{1}^{\sigma}\\ J_{1}^{\sigma}\end{pmatrix}=\begin{pmatrix}L_{eV}^{\sigma}&L_{e\theta}^{\sigma}\\ L_{hV}^{\sigma}&L_{h\theta}^{\sigma}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (5)
@@
where the spin polarized Onsager matrix elements $L_{eV}^{\sigma}$, $L_{e\theta}^{\sigma}$, $L_{hV}^{\sigma}$ and $L_{h\theta}^{\sigma}$ are derived in Eq. (65) of SM. Using Eq. (5), the spin polarized transport coefficients such as Conductance ($G^{\sigma}$), Seebeck Coefficient ($S^{\sigma}$), Peltier Coefficient ($\Pi^{\sigma}$), and thermal conductance ($K^{\sigma}$) are defined as,
$G^{\sigma}$ $=L_{eV}^{\sigma},\quad S^{\sigma}=\frac{L_{e\theta}^{\sigma}}{L_{eV}^{\sigma}},\quad\Pi^{\sigma}=\frac{L_{hV}^{\sigma}}{L_{eV}^{\sigma}},\quad K^{\sigma}=L_{h\theta}^{\sigma}-\frac{L_{hV}^{\sigma}L_{e\theta}^{\sigma}}{L_{eV}^{\sigma}}.$ (6)
@@
The output power in terminal 1, $P=I_{1}V$ is maximum at a voltage bias $V=L_{e\theta}\tau/4L_{eV}$ and is given as $P_{max}=L_{e\theta}^{2}\tau^{2}/4L_{eV}$, see Sec. II.1.2 for QH and III A 2 for QSH in SM. The efficiency at maximum power, i.e., $\eta|_{P_{max}}=\frac{P_{max}}{J}=\frac{\eta_{c}}{2}\frac{Z\theta}{Z\theta+2}$ (see, Eq. (34) for QH and Eq. (75) for QSH in SM), $J=J_{1}$ is the heat current from the hotter terminal 1. $Z\theta=\frac{GS^{2}\theta}{K}$ is the figure of merit. Similarly, maximum efficiency, i.e., $\eta_{max}=\eta_{c}\frac{\sqrt{Z\theta+1}-1}{\sqrt{Z\theta+1}+1}$ can be derived (see Eq. (38) for QH and Eq. (79) for QSH in SM).
@@
In Fig. 2(a, b), we parametrically plot the efficiency ($\eta/\eta_{c}$) and the power ($P/P_{max}$) in linear response regime for both QH and QSH QHE, calculated using Eqs. (3-6) and observe that it achieves the Curzon-Ahlborn limit at $V=0.057k_{B}\theta/e$ and further increasing $V$, power decreases and efficiency increases (see blue (red) dashed curve for $V=0.081k_{B}\theta/e$ and blue (red) dotted curve for $V=0.096k_{B}\theta/e$) . We considered threshold energies of QPC 1 and QPC 2 to be $E_{1}^{{}^{\prime}}=E_{1}+eV_{g}$ and $E_{2}^{{}^{\prime}}=E_{2}+eV_{g}$ respectively, with $E_{1}=E_{2}$. If $E_{1}\neq E_{2}$; the conclusions don’t change, see, Secs. II.1.4 (II.2.4) for QH and III A 3 (III B 4) for QSH for the linear (non-linear) transport in SM.
@@
In Ref. [7], it has been reported that a QH setup cannot work as a QR due to time-reversal symmetry breaking (TRS). Conversely, the same setup with QSH edge modes can act as both QHE and QR because TRS is preserved [12]. In the QH setup as in [7], the asymmetric parameter (AP), i.e., the Seebeck to Peltier coefficient ratio, is zero or infinity, affecting its refrigeration capability [31, 12]. In the QSH setup, AP is always finite since TRS is preserved [12]. However, this is not true for all QH setups. The QH setup in Fig. 1(a) and the QSH setups in Fig. 1(b) can act as both QHE and QR because their AP is finite (see Sec. II.1.4 of SM).
@@
For QR, we can calculate the cooling power at the maximum coefficient of performance ($\textbf{J}|_{\eta^{r}_{max}}$) and the maximum coefficient of performance ($\eta_{max}^{r}$) for both QH and QSH setups. The coefficient of performance (COP = $\textbf{J}/P$) is generally defined as the ratio of heat extracted from the cooler terminal ($\textbf{J}=-J_{2}$) to the power absorbed ($P$) by the setup. The maximum COP ($\eta^{r}_{max}$) is achieved using the condition $\frac{d(COP)}{dV}=0$ and is given as $\eta^{r}_{max}=\eta^{r}_{c}\frac{\sqrt{Z\theta+1}-1}{\sqrt{Z\theta+1}+1}$, where $\eta^{r}_{c}$ is the Carnot coefficient of performance. As shown in Fig. 2(c), $\eta^{r}_{max}$ reaches $0.97\eta_{c}^{r}$ at $\textbf{J}|_{\eta^{r}_{max}}$ at $V=0.115k_{B}\theta/e$. Similar to the QH setup, we parametrically plot the cooling power J vs coefficient of performance $\eta^{r}$ in Fig. 2(d) for QSH setup.
@@
As shown in Sec. II.1.2, for QH and Sec. III.1.2 for QSH setups in the SM, both $\eta_{max}$ and $\eta|_{P_{max}}$ rely on $Z\theta$ for a QHE. Similarly, for QR, $\eta^{r}_{max}$ depend on $Z\theta$. To achieve Curzon-Ahlborn efficiency ($\frac{\eta_{c}}{2}$), $Z\theta$ must significantly exceed 1. The key lies in having a large Seebeck coefficient ($S$) and a small thermal conductance ($K$), both occurring in the same configuration. In both of our QH and QSH setups, we find that the Seebeck coefficient is notably large, while the thermal conductivity is remarkably small..

### v3

# Reaching Van den Broeck limit in linear response and Whitney limit in nonlinear response in edge mode quantum thermoelectrics and refrigeration

@@ Abstract
Quantum heat engines and quantum refrigerators are proposed in three-terminal quantum Hall (QH) and quantum spin Hall (QSH) setups with a voltage-temperature probe in both the linear and nonlinear transport regimes. In the linear response regime, we find that efficiency at maximum power approaches the Van den Broeck limit in both QH and QSH setups. Similarly, in nonlinear response, we find that efficiency at maximum power reaches the Whitney bounds. This is for the first time, we see that in the same setup and using quantum point contacts, the thermoelectric efficiency limits in linear and nonlinear response being achieved.
@@
Introduction: The quantum Hall resistance [1] in two-dimensional electron gas in the presence of a normal uniform magnetic field at low temperature arises as a consequence of chiral edge modes [2], which is famously known as quantum Hall effect. These edge modes are robust against disorder and are immune to backscattering [3] being topologically protected. Complimentary to chiral edge modes, helical edge modes are seen in quantum spin Hall experiments with Mercury Telluride/Cadmium Telluride heterostructures [4, 5, 6], which are also topologically protected. The study of thermoelectric properties in quantum Hall (QH) and quantum spin Hall (QSH) setups has received considerable attention, particularly through the application of the Landauer-Büttiker formalism [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This formalism forms the basis for understanding the behavior of quantum heat engines and quantum refrigerators in time-independent steady-state quantum thermoelectric studies, which is the main focus of our investigation. Similarly, in the realm of quantum thermodynamics, heat engines and refrigerators are also broadly categorized into finite-time cyclic systems [25, 26, 27, 28] and autonomous (steady-state) systems [29, 30, 31]. Quantum transport can be described in two regimes. First is linear, wherein the applied voltage and temperature biases are quite small, and thus leads to a linear relation between current and the temperature or voltage biases applied [32]. The second is nonlinear, wherein the applied voltage and temperature biases are not small. In the linear response regime, the efficiency at maximum power ($\eta|_{P_{max}}$) [33] and the maximum efficiency ($\eta_{max}$) are closely related to the figure of merit $ZT\hskip 2.84544pt\left(=\frac{GS^{2}T}{K}\right)$, where $G$ = conductance, $S$ = Seebeck coefficient, $K$ = thermal conductance and $T$ is the reference temperature of the setup. In this regime, there exists a bound to $\eta|_{P_{max}}$, which is given by Van den Broeck (VDB) [34] limit with the efficiency ($\eta|_{P_{max}}$)= $\eta_{c}/2$, where $\eta_{c}=1-T_{2}/T_{1}$ is the Carnot efficiency with $T_{1}$ and $T_{2}$ being the temperatures of the hot and cold reservoirs respectively [32]. To achieve this, it is necessary to have a large $ZT$. This can be accomplished by ensuring a high Seebeck coefficient ($S$) and low thermal conductance ($K$). Similarly, in the non-linear transport regime, there also exists an upper bound to $\eta|_{P_{max}}$, the Whitney limit [13, 14], i.e., efficiency at maximum power output ($\eta|_{P_{max}^{Wh}}$) = $\eta|_{P_{max}^{Wh}}$ = $\eta_{c}/(1+0.936(1+T_{2}/T_{1}))$. This is the nonlinear equivalent of the VDB efficiency in linear response and has been derived utilizing BoxCar-type transmission. The study of two-terminal setups in nonlinear response was extended to a three-terminal setup with a voltage probe, and it was shown that the Whitney limits are universal [35]. Similarly, in [36], it has been proved that a generic mesoscopic thermoelectric heat engine can be highly efficient only when the transmission is Box-car type. Boxcar-type scattering can be implemented via a chain of quantum dots [13, 14, 35]. However, electron-electron interaction between quantum dots, which can lead to coulomb blockade effects, is neglected in the calculation in Refs. [13, 14, 35]. This assumption is quite extreme, as in real systems, having a chain of quantum dots without the coulomb blockade effect is almost impossible to implement [32]. In contrast, this work considers quantum point contact (QPC) type scattering, which is easier to implement experimentally [37, 38, 39, 40, 41] than the boxcar-type transmission. In this work, we reach the Whitney bound in the nonlinear transport regime using QPC-type scattering. Previously, in Refs. [13, 14, 35] a QH setup was only considered; however, in this letter, we consider both QH and QSH setups, albeit with the easier-to-implement QPC type contacts for both linear as well as non-linear response regimes.
@@
This letter is divided into two parts. In the first part, we investigate the thermoelectric transport of the QH and QSH setups within the linear response regime and show that efficiency at maximum power approaches Van den Broeck for both QH and QSH setups. In the second part, we focus on non-linear transport, wherein we see that the efficiency at maximum power approaches the Whitney limit for both QH and QSH setups. The technical details of the calculation are presented in the Supplementary Material (SM).
Thermoelectricity in QH and QSH setups: We start by discussing the general theory for thermoelectric transport in QH as well as QSH setups within the Landaeur-Buttiker scattering theory in both linear and non-linear regimes [32, 7]. The net charge and heat currents in a multiterminal QH sample, see Sec. I of SM, are given as,
@@
where the transmission probability for an electron to scatter from terminal $\beta$ to terminal $\alpha$ is $\mathcal{T}_{\alpha\beta}$. Here, $f_{\alpha(\beta)}=\left(1+e^{(E-\mu_{\alpha(\beta)})/k_{B}T_{\alpha(\beta)}}\right)^{-1}$, $\mu_{\alpha(\beta)}$ and $T_{\alpha(\beta)}$ are the Fermi function, chemical potential and temperature of terminal $\alpha(\beta)$ respectively.
@@
Similarly, for QSH setup, with helical edge modes, the currents carry an extra spin degree of freedom and are written as $I_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{\alpha}^{\sigma}$ and $J_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{\alpha}^{\sigma}$, where [15, 23],
@@
$\begin{split}I_{\alpha}^{\sigma}&=\frac{e}{h}\int_{-\infty}^{\infty}dE\sum_{\beta}\mathcal{T}_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\\ J_{\alpha}^{\sigma}&=\frac{1}{h}\int_{-\infty}^{\infty}dE(E-\mu_{\alpha})\sum_{\beta}\mathcal{T}_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\end{split}$ (2)
@@
with the transmission probability for an electron to transmit from terminal $\beta$ with spin $\rho$ to terminal $\alpha$ with spin $\sigma$ being $\mathcal{T}_{\alpha\beta}^{\sigma}=\sum_{\rho\in\{\uparrow,\downarrow\}}\mathcal{T}_{\alpha\beta}^{\sigma\rho}$. In this three-terminal setup (both QH and QSH) as shown in Fig. 1, $T_{1}=T+\tau_{1},T_{2}=T+\tau_{2}$ and $T_{3}=T+\tau_{3}$, where $T$ is the reference temperature and $\tau_{1},\tau_{2}$ and $\tau_{3}$ are the temperature biases applied in terminals 1, 2 and 3 respectively. For our setups, we apply temperature bias only at terminal 1, i.e., $\tau_{1}=\tau$ and $\tau_{2}=0$. We study the thermoelectric properties for a QHE, considering terminal 3 to be a VT probe. We also assume both terminals 1 and 2 to be current probes with $V_{1}=-V$ and $V_{2}=0$.
Quantum point contact vs. boxcar-type transmission: In this section, we discuss the two types of transmission, which we have extensively considered in this work. The first is the quantum point contact (QPC)-type tunneling [44] and the other one is the boxcar-type [13, 14].
A quantum point contact can be modeled by a potential
$V(x,y)=V_{0}-\frac{1}{2}m\omega_{x}^{2}x^{2}+\frac{1}{2}m\omega_{y}^{2}y^{2},$ (3)
where $m$ is the effective mass of the electron and the electron feels the potential $V(x,y)$ when the QPC is present in a setup. As derived in Ref. [44], the transmission probability ($T_{QPC}$) for this constriction is given as
$T_{QPC}=\sum_{n}\frac{1}{1+\text{exp}\left[-2\pi(E-E_{n})/\hbar\omega_{x}\right]},$ (4)
where, $E_{n}=V_{0}+\hbar\omega_{y}\left(n+1/2\right)$, with $n=0,1,2,...$. We consider $E_{n}$ in the 1D limit, where only a few subbands are below the Fermi energy $E_{F}$. We assume the length along $x$-direction ($l_{x}$) to be much larger than that of $y$-direction ($l_{y}$), therefore only $E_{n}$ is quantized along $y$-direction. one can consider $n=1$ and $\omega_{x},\omega_{y}=\omega$ and we have the transmission probability ($T_{QPC}$) as,
$T_{QPC}=\frac{1}{1+e^{-2\pi(E-E_{1})/\hbar\omega}},$ (5)
which has been shown in Fig. 2(a), where below a certain energy value $E_{1}$, it vanishes, and above this, it is always one. Here, $\hbar\omega$ is the width of the step. In our work, we have capacitively connected a gate voltage $V_{g}$ to the QPC-type constriction, giving $V(x,y)\implies V(x,y)+eV_{g}$, which implies $E_{1}\rightarrow E_{1}+eV_{g}$. We denote $E_{1}+eV_{g}$ as $E_{1}^{\prime\prime}$.
Figure 2: (a) QPC type, (b) Boxcar-type transmission.
Similarly, the boxcar-type of transmission is 1 in the energy range $E_{1}<E<E_{1}^{\prime}$, otherwise zero [13, 14], See Fig. 2(b). This transmission function will be used while we discuss Whitney’s approach to studying the thermoelectrics of a two-terminal QH and QSH setup. As discussed by Whitney, one can generate a boxcar-type of transmission by considering many quantum dots in a region between two terminals [13, 14]. When there is only a single quantum dot with a single energy level, the transmission probability via it will be the resonant-tunneling type [45]. With one quantum dot, the electron will hop into it from one terminal and later hop out into the other terminal. But when one increases the number of quantum dots, each having a single energy level, then the electrons hop repeatedly between the quantum dots while going from one terminal to the other, which gives a boxcar-type transmission, which is $1$ in between two arbitrary energy levels $E_{1}$ and $E_{1}^{\prime}$ as shown in Fig. 2(b). When $E_{1}^{\prime}\rightarrow\infty$, the boxcar type transmission becomes QPC-like, always 1 above $E_{1}$ and zero otherwise.
However, as discussed in the introduction, implementing the boxcar-type transmission is more difficult than implementing the QPC-type transmission. In fact, QPC was the first type of tunneling used to study the quantum transport in 2DEGs [37, 38, 39, 40] such as GaAs-AlGaAs heterostructure. Most of the ballistic or edge mode transport can be done much more conveniently with the help of QPC in mesoscopic samples. For the implementation of QPC, one needs to control electrostatically the local electrostatic potential, which is felt by the electron with the help of metallic split gates [46, 47, 48]. Here, in this work, with the help of only the QPC-type of tunneling, one can study the nonlinear transport, which can help reach the Whitney bounds.
@@
Linear transport in QH and QSH setups: In the linear response regime for the QH case, i.e., when $\mu_{\alpha}(=eV_{\alpha})\ll k_{B}T$ and ($T_{\alpha}-T)\ll T$, one can see that the current (both charge and heat) varies linearly with the voltage and temperature biases applied across the sample [32]. Since, terminal 3 is a VT probe i.e., $I_{3}=J_{3}=0$, the charge and heat currents in terminal 1 can be written as,
$\begin{pmatrix}I_{1}\\ J_{1}\end{pmatrix}=\begin{pmatrix}L_{eV}&L_{eT}\\ L_{hV}&L_{hT}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (6)
@@
where the Onsager matrix elements are $L_{eV},L_{eT},L_{hV},L_{hT}$, see Eq. (31) in SM. These Onsager matrix elements are used to find the transport coefficients such as conductance ($G$), Seebeck coefficient ($S$), Peltier coefficient ($\Pi$), and the thermal conductance ($K$) and are given as,
$G$ $=L_{eV},\quad S=\frac{L_{eT}}{L_{eV}},\quad\Pi=\frac{L_{hV}}{L_{eV}},\quad K=L_{hT}-\frac{L_{hV}L_{eT}}{L_{eV}}.$ (7)
@@
Using a similar procedure, one can also derive the Onsager matrix elements in the QSH case for charge current $I_{\alpha}^{\sigma}$ and heat current $J_{\alpha}^{\sigma}$ by imposing the VT probe condition ($I_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{3}^{\sigma}=0$ and $J_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{3}^{\sigma}=0$) and they can be written in the same form as Eq. (6) (see, Sec. II.1.1 in SM as
$\begin{pmatrix}I_{1}^{\sigma}\\ J_{1}^{\sigma}\end{pmatrix}=\begin{pmatrix}L_{eV}^{\sigma}&L_{eT}^{\sigma}\\ L_{hV}^{\sigma}&L_{hT}^{\sigma}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (8)
@@
where the spin polarized Onsager matrix elements $L_{eV}^{\sigma}$, $L_{eT}^{\sigma}$, $L_{hV}^{\sigma}$ and $L_{hT}^{\sigma}$ are derived in Eq. (81) of SM. Using Eq. (8), the spin polarized transport coefficients such as Conductance ($G^{\sigma}$), Seebeck Coefficient ($S^{\sigma}$), Peltier Coefficient ($\Pi^{\sigma}$), and thermal conductance ($K^{\sigma}$) are defined as,
$G^{\sigma}$ $=L_{eV}^{\sigma},\quad S^{\sigma}=\frac{L_{eT}^{\sigma}}{L_{eV}^{\sigma}},\quad\Pi^{\sigma}=\frac{L_{hV}^{\sigma}}{L_{eV}^{\sigma}},\quad K^{\sigma}=L_{hT}^{\sigma}-\frac{L_{hV}^{\sigma}L_{eT}^{\sigma}}{L_{eV}^{\sigma}}.$ (9)
@@
The output power in terminal 1, $P=I_{1}V$ is maximum at a voltage bias $V=L_{eT}\tau/4L_{eV}$ and is given as $P_{max}=L_{eT}^{2}\tau^{2}/4L_{eV}$, see Sec. I.1.2 for QH and II.1.2 for QSH in SM. The efficiency at maximum power, i.e., $\eta|_{P_{max}}=\frac{P_{max}}{J}=\frac{\eta_{c}}{2}\frac{ZT}{ZT+2}$ (see, Eq. (34) for QH and Eq. (85) for QSH in SM, $J=J_{1}$ is the heat current from the hotter terminal 1. $ZT=\frac{GS^{2}T}{K}$ is the figure of merit. Similarly, maximum efficiency, i.e., $\eta_{max}=\eta_{c}\frac{\sqrt{ZT+1}-1}{\sqrt{ZT+1}+1}$ can be derived (see Eq. (40) for QH and Eq. (89) for QSH in SM).
@@
In Fig. 3(a, b), we parametrically plot the efficiency ($\eta/\eta_{c}$) and the power ($P/P_{max}$) in linear response regime for both QH and QSH QHE, calculated using Eqs. (6-9) and observe that it achieves the Van den Broeck limit at $V=0.057k_{B}T/e$ and further increasing $V$, power decreases and efficiency increases (see blue (red) dashed curve for $V=0.081k_{B}T/e$ and blue (red) dotted curve for $V=0.096k_{B}T/e$) . We considered threshold energies of QPC 1 and QPC 2 to be $E_{1}^{\prime\prime}=E_{1}+eV_{g}$ and $E_{2}^{\prime\prime}=E_{2}+eV_{g}$ respectively, with $E_{1}=E_{2}$. If $E_{1}\neq E_{2}$; the conclusions don’t change, see, Secs. I.1.4 (I.2.4) for QH and II.1.3 (I.2.4) for QSH for the linear (non-linear) transport in SM.
@@
In Ref. [9], it has been reported that a QH setup cannot work as a QR due to time-reversal symmetry breaking (TRS). Conversely, the same setup with QSH edge modes can act as both QHE and QR because TRS is preserved [15]. In the QH setup as in [9], the asymmetric parameter (AP), i.e., the Seebeck to Peltier coefficient ratio, is zero or infinity, affecting its refrigeration capability [49, 15]. In the QSH setup, AP is always finite since TRS is preserved [15]. However, this is not true for all QH setups. The QH setup in Fig. 3(a) and the QSH setups in Fig. 3(b) can act as both QHE and QR because their AP is finite (see Sec. I.1.4 of SM).
@@
For QR, we can calculate the cooling power at the maximum coefficient of performance ($\textbf{J}|_{\eta^{r}_{max}}$) and the maximum coefficient of performance ($\eta_{max}^{r}$) for both QH and QSH setups. The coefficient of performance (COP = $\textbf{J}/P$) is generally defined as the ratio of heat extracted from the cooler terminal ($\textbf{J}=-J_{2}$) to the power absorbed ($P$) by the setup. The maximum COP ($\eta^{r}_{max}$) is achieved using the condition $\frac{d(COP)}{dV}=0$ and is given as $\eta^{r}_{max}=\eta^{r}_{c}\frac{\sqrt{ZT+1}-1}{\sqrt{ZT+1}+1}$, where $\eta^{r}_{c}$ is the Carnot coefficient of performance. As shown in Fig. 3(c), $\eta^{r}_{max}$ reaches $0.97\eta_{c}^{r}$ at $\textbf{J}|_{\eta^{r}_{max}}$ at $V=0.115k_{B}T/e$. Similar to the QH setup, we parametrically plot the cooling power J vs coefficient of performance $\eta^{r}$ in Fig. 3(d) for QSH setup.
@@
As shown in Sec. I.1.2, for QH and Sec. II.1.2 for QSH setups in the SM both $\eta_{max}$ and $\eta|_{P_{max}}$ rely on $ZT$ for a QHE. Similarly, for QR, $\eta^{r}_{max}$ depend on $ZT$. To achieve Van den Broeck efficiency ($\frac{\eta_{c}}{2}$), $ZT$ must significantly exceed 1. The key lies in having a large Seebeck coefficient ($S$) and a small thermal conductance ($K$), both occurring in the same configuration. In both of our QH and QSH setups, we find that the Seebeck coefficient is notably large, while the thermal conductivity is remarkably small.